OpenStudy (anonymous):

McPherson Cement sells cement for $65 plus $22 for each yard of cement. How many yards of cement can be purchased for $417?

OpenStudy (anonymous):

65+22x≤417 22x≤352 x≤16 so 16 yards

OpenStudy (lilg132):

T = c + 22(y) t = total c = cement price y = yards 417 = 65 + 22(y) 417 - 65 = 22y 352 = 22y y = 16 16 yards